In a large city, two newspapers are published, the Sun and the Post. The circulation departments report that 22% of the city’s households have a subscription to the Sun and 35% subscribe to the Post.

1) What is the probability of a household subscribing to both the Sun and the Post?

2) What is the probability of a household subscribing to either the Sun or the Post?

PLEASE SHOW WORK! I'd like to see how to solve it so I can do this for future problems!
Answer #1

(a)

Let

A = subscription to Sun

B = subscription to Post

Given:
P(A) = 0.22

P(B) = 0.35

By Multiplication Theorem:

PLA B) = P(A) × P(B) = 0.22 × 0.35 = 0.077

So,

P(Both Sun & the Post) = 0.077

(b)
By Addition Theorem:

Pl A | B) = P(A) + P(B) _ P(A0 B) = 0.22 + 0.35-0.077 = 0.493

So,

P(either the Sun or the Post) = 0.493